a 30 year-old father has 4 year-old son. In how many years will the father be thrice as old as his son ? Show algebraic solution
Answer by aliafzal14(13)

hello
first we take a difference of their age
let father's age = F
and Son's age = S

F - S = 26 eq-1 (30-4=26)
After X years :
F = 3S put in eq-1
3S - S = 26
2S = 26
S = 13 so F = 3S = 39
X = 9 years

so it is shown that after 9 years, father will be thrice as old as his son